OpenStudy (anonymous):

3x + 5y = 78 2x - y = 0 The point of intersection of the lines has an x-coordinate of ___. @salehgeek

OpenStudy (anonymous):

you can rewrite the second equation as \(2x=y\) and then replace the \(y\) in the first equation by \(2x\)

OpenStudy (anonymous):

x=6 and y= 12

OpenStudy (anonymous):

to get \(3x+5(2x)=78\) and then solve for \(x\)

OpenStudy (anonymous):

@brittnicoleee_ do you have the answer you are looking for?

OpenStudy (anonymous):

yes x + 5y = -2 2x + y = 5 The point of intersection of the lines has a y-coordinate of ___.

OpenStudy (anonymous):

If not I can help.

OpenStudy (anonymous):

Ok just making sure.

OpenStudy (anonymous):

x + 5y = -2 2x + y = 5 The point of intersection of the lines has a y-coordinate of ___.